OpenStudy (thatonegirl_):

Can someone tell me what I'm doing wrong? Cotθ=-2/5. π/2< θ<π. Find the other five trigonometric functions of θ.

OpenStudy (thatonegirl_):

Tanθ=-5/2, so isn't cosθ=-2 and sinθ=5?

OpenStudy (anonymous):



OpenStudy (anonymous):

find \(h\) via pythagoras then you can take any trig ratio you like

OpenStudy (anonymous):

also it is impossible for sine to be 5 sine is never bigger than one it is ratios you need

OpenStudy (campbell_st):

so its 2nd quadrant 

OpenStudy (thatonegirl_):

Right.. But the tangent identity is sinθ/cosθ

OpenStudy (anonymous):

just because \[\cot(x)=\frac{\cos(x)}{\sin(x)}\] doesnt mean cosine is the numerator and sine is the denominator

OpenStudy (thatonegirl_):

oh .. -.-

OpenStudy (thatonegirl_):

ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@campbell_st picture is better than mine, it shows you what quadrant you are in i usually just ignore that, and would figure out the signs of the trig functions by being told you are in quadrant 2

OpenStudy (thatonegirl_):

yeah i got that part, thank you :D